# Generalizable, Fast, and Accurate DeepQSPR with fastprop

### v1

@@ Abstract
Quantitative Structure-Property/Activity Relationship studies, often referred to interchangeably as QS(P/A)R, seek to establish a mapping between molecular structure and an arbitrary Quantity of Interest (QOI). Since its inception this was done on a QOI-by-QOI basis with new descriptors being devised by researchers to specifically map to their QOI. This continued for years and culminated in packages like DRAGON (later E-dragon), PaDEL-descriptor (and padelpy), Mordred, and many others. The sheer number of different packages resulted in the creation of ‘meta-packages’ which served only to aggregate these other calculators, including tools like molfeat, ChemDes, Parameter Client, and AIMSim. Despite the creation of these aggregation tools, QSPR studies have continued to focus on QOI-by-QOI studies rather than attempting to create a generalizable approach which is capable of modeling across chemical domains.
One contributing factor to this is that QSPR studies have relied almost exclusively on linear methods for regression. Efforts to incorporate Deep Learning (DL) as a regression technique (Deep-QSPR), which would be capable of capturing the non-linear behavior of arbitrary QOIs, have instead focused on using molecular fingerprints as inputs. The combination of bulk molecular-level descriptors with DL has remained largely unexplored, in significant part due to the orthogonal domain expertise required to combine the two. Generalizable QSPR has turned to learned representations primarily via message passing graph neural networks. This approach has proved remarkably effective but is not without drawbacks. Learning a representation requires large datasets to avoid over-fitting or even learn at all, loses interpretability since an embedding’s meaning can only be induced retroactively, and needs significant execution time given the complexity of the underlying message passing algorithm. This paper introduces fastprop, a software package and general Deep-QSPR framework that combines a cogent set of molecular descriptors with DL to achieve state-of-the-art performance on datasets ranging from tens to tens of thousands of molecules. fastprop is designed with Research Software Engineering best practices and is free and open source, hosted at github.com/jacksonburns/fastprop.
@@ Introduction
While iterations on LRs and novel approaches to low-data regimes have been in development, the classical QSPR community has continued their work. A turning point in this domain was the release of mordred, a fast and well-developed package capable of calculating more than 1600 molecular descriptors. Critically this package was fully open source and written in Python, allowing it to readily interoperate with the world-class Python DL software ecosystem that greatly benefitted the LR community. Now despite previous evidence that molecular descriptors cannot achieve generalizable QSPR in combination with DL, the opposite is shown.
@@
At its core the fastprop ‘architecture’ is simply the mordred molecular descriptor calculator 222The original mordred package is no longer maintained. fastprop uses a fork of mordred called mordredcommunity that is maintained by community-contributed patches (see github.com/JacksonBurns/mordred-community). Multiple descriptor calculators from the very thorough review by McGibbon et al. [16] could be used instead, though none are as readily interoperable as mordred. [17] connected to a Feedforward Neural Network (FNN) implemented in PyTorch Lightning [18] (Figure 1). The user simply specifies a set of SMILES [19], a linear textual encoding of molecules, and their corresponding properties. fastprop automatically calculates and caches the corresponding molecular descriptors with mordred, re-scales both the descriptors and the targets appropriately, and then trains an FNN with to predict the indicated target properties. By default this FNN is two hidden layers with 1800 neurons each connected by ReLU activation functions, though the configuration can be readily changed via the command line interface or configuration file. fastprop owes its success to the cogent set of descriptors assembled by the developers of mordred, the ease of training FNNs with modern software like PyTorch Lightning, and the careful application of Research Software Engineering best practices that make it as user friendly as the best-maintained alternatives.
@@
# pah.yml # generic args output_directory: pah random_seed: 55 problem_type: regression # featurization input_file: pah/arockiaraj_pah_data.csv target_columns: log_p smiles_column: smiles descriptors: all # preprocessing zero_variance_drop: False colinear_drop: False # training number_repeats: 4 number_epochs: 100 batch_size: 64 patience: 15 train_size: 0.8 val_size: 0.1 test_size: 0.1 sampler: random
@@
There are a number of established molecular property prediction benchmarks commonly used in LR studies, especially those standardized by MoleculeNet [24]. Principal among them are QM8 [25] and QM9 [26], often regarded as the standard benchmark for property prediction. These are important benchmarks and are included here for completeness, though the enormous size and rich coverage of chemical space (inherent in the design of the combinatorial datasets) means that nearly all model architectures are highly accurate.
Real world datasets, particularly those common in QSPR studies, often number in the hundreds. To demonstrate the applicability of fastprop to these regimes, many smaller datasets are selected including some from the QSPR literature that are not established benchmarks. These studies relied on more complex and slow modeling techniques (ARA) or the design of a bespoke descriptor (PAHs) and have not yet come to rely on learned representations as a go-to tool. In these data-limited regimes where LRs sometimes struggle, fastprop and its intuition-loaded initialization are highly powerful. To emphasize this point further, the benchmarks are presented in order of size, descending, for first regression and then classification tasks. Consistently fastprop is able to match the performance of LRs on large datasets (O(10,000)+ entries) and compete with or exceed their performance on small datasets (O(1,000) entries).
@@
All benchmarks use 80% of the dataset for training (selected randomly, unless otherwise stated), 10% for validation, and holdout the remaining 10% for testing (unless otherwise stated). Sampling is performed using the astartes package [27] which implements a variety of sampling algorithms and is highly reproducible.
Results for fastprop are reported as the average value of a metric and its standard deviation across a number of repetitions (repeated re-sampling of the dataset). The number of repetitions is chosen to either match referenced literature studies or else increased from two until the performance no longer meaningfully changes. Note that this is not the same as cross-validation.
For performance metrics retrieved from literature it is assumed that the authors optimized their respective models to achieve the best possible results; therefore, fastprop metrics are reported after model optimization using the fastprop train ... --optimize option. When results are generated for this study using Chemprop, the default settings are used except that the number of epochs is increased to allow the model to converge and batch size is increased to match dataset size and speed up training.
When reported, execution time is as given by the unix time command using Chemprop version 1.6.1 on Python 3.8 and includes the complete invocation of Chemprop, i.e. time chemprop_train .... The insignificant time spent manually collating Chemprop results (Chemprop does not natively support repetitions) is excluded. fastprop is run on version 1.0.0b2 using Python 3.11 and timing values are reported according to its internal time measurement which was verified to be nearly identical to the Unix time command. The coarse comparison of the two packages is intended to emphasize the scaling of LRs and Deep-QSPR and that fastprop is, generally speaking, much faster. All models trained for this study were run on a Dell Precision series laptop with an NVIDIA Quadro RTX 4000 GPU and Intel Xeon E-2286M CPU.
@@
Regression Datasets
See Table LABEL:regression_results_table for a summary of all the regression dataset results. Especially noteworthy is the performance on the ESOL and PAH datasets, which dramatically surpass literature best and have only 55 datapoints, respectively. Citations for the datasets themselves are included in the sub-sections of this section.
@@
QM9 | ~134 | MAE | 0.0047a | 0.0069 | 0.0081a |
OCELOTv1 | ~25 | MAE | 0.128b | 0.158 | 0.140b |
QM8 | ~22 | MAE | 0.016a | 0.018 | 0.019a |
ESOL | ~1.1 | RMSE | 0.55c | 0.64 | 0.67c |
FreeSolv | ~0.6 | RMSE | 0.82c | 1.33 | 1.26c |
Flash | ~0.6 | RMSE | 13.2e | 13.3 | 21.2* |
YSI | ~0.4 | MAE | 22.3f | 13.6 | 28.9* |
HOPV15g | ~0.3 | MAE | 1.32g | 1.55 | 1.60 |
Fubrain | ~0.3 | RMSE | 0.44h/0.83d | 0.19/0.74 | 0.22*/0.97d |
PAH | ~0.06 | R2 | 0.96i | 0.98 | 0.59* |
a [9] b [10] (summary result is geometric mean across tasks) c [8] d [14] (delta-learning instead of direct prediction) e [28] f [29] g [30] (uses a subset of the complete HOPV15 dataset) h [22] i [4] * These results were generated for this study.
@@
fastprop achieves 0.0069 0.0002 mean absolute error, whereas Chemprop achieves 0.00814 0.00001 and the UniMol framework manages 0.00467 0.00004. This places the fastprop framework ahead of previous learned representation approaches but still trailing UniMol. This is not completely unexpected since UniMol encodes 3D information from the dataset whereas Chemprop and fastprop use only 2D. Future work could evaluate the use of 3D-based descriptors to improve fastprop performance in the same manner that UniMol has with LRs.
OCELOTv1
The Organic Crystals in Electronic and Light-Oriented Technologies (OCELOTv1) dataset, originally described by Bhat et al. [31], maps 15 quantum mechanics descriptors and optoelectronic properties to ~25k chromophoric small molecules. The literature best model is the Molecular Hypergraph Neural Network (MHNN) [10] which specializes in learned representations for optoelectronic properties and also includes Chemprop as a baseline for comparison. They used a 70/10/20 random split with three repetitions and final performance reported is the average across those three repetitions.
As done in the reference study, the MAE for each task is shown in Table LABEL:ocelot_results_table. Meanings for each abbreviation are the same as in the original database publication [31]. The geometric mean across all tasks, which accounts for the different scales of the target values better than the arithmetic mean, is also included as a summary statistic. Note also that the relative percentage performance difference between fastprop and chemprop (fast/chem) and fastprop and MHNN (fast/MHNN) are also included.
Target | fastprop | Chemprop | fast/chem | MHNN | fast/mhnn |
HOMO | 0.328 | 0.330 | -0.7% | 0.306 | 7.1% |
LUMO | 0.285 | 0.289 | -1.5% | 0.258 | 10.4% |
H-L | 0.555 | 0.548 | 1.4% | 0.519 | 7.0% |
VIE | 0.210 | 0.191 | 9.9% | 0.178 | 18.0% |
AIE | 0.201 | 0.173 | 16.2% | 0.162 | 24.0% |
CR1 | 0.057 | 0.055 | 3.3% | 0.053 | 7.2% |
CR2 | 0.057 | 0.053 | 7.2% | 0.052 | 9.3% |
HR | 0.108 | 0.133 | -18.5% | 0.099 | 9.5% |
VEA | 0.194 | 0.157 | 23.5% | 0.138 | 40.5% |
AEA | 0.188 | 0.154 | 21.8% | 0.124 | 51.3% |
AR1 | 0.055 | 0.051 | 8.2% | 0.050 | 10.4% |
AR2 | 0.049 | 0.052 | -5.1% | 0.046 | 7.3% |
ER | 0.100 | 0.098 | 1.9% | 0.092 | 8.5% |
S0S1 | 0.284 | 0.249 | 14.3% | 0.241 | 18.0% |
S0T1 | 0.222 | 0.150 | 48.3% | 0.145 | 53.4% |
G-Mean | 0.151 | 0.140 | 7.7% | 0.128 | 17.9% |
fastprop ’trades places with Chemprop, outperforming on four of the metrics (LUMO, HR, AR2) and under-performing on others. Overall the geometric mean of MAE across all the tasks is ~8% higher, though this result may not be statistically significant. Both fastprop and Chemprop are outperformed by the bespoke MHNN model, which is not itself evaluated on any other common property prediction benchmarks.
Although fastprop is not able to reach state-of-the-art accuracy on this dataset this result is still promising. None of the descriptors implemented in mordred were designed to specifically correlate to these QM-derived targets, yet the FNN is able to learn a representation which is nearly as informative as Chemprop. The fact that a bespoke modeling approach is the most performant is not surprising and instead demonstrates the continued importance of expert input on certain domains. Were some of this expertise to be oriented toward the descriptor generation software mordredcommunity, new descriptors could be added to address this apparent shortcoming.
QM8
Quantum Machine 8 (QM8) is the predecessor to QM9 first described in 2015 [25]. It follows the same generation procedure as QM9 but includes only up to eight heavy atoms for a total of approximately 22k molecules. Again, this study used the dataset as prepared by MoleculeNet [24] and compares to the UniMol [9] set of benchmarks as a reference point, wherein they used the same data splitting procedure described previously but regressed all 12 targets in QM8.
UniMol achieved an average MAE across all tasks of 0.00156 0.0001, fastprop approaches that performance with 0.0178 0.003, and Chemprop trails both frameworks with 0.0190 0.0001. Much like with QM9 fastprop outperforms LR frameworks until 3D information is encoded with UniMol. As previously stated this is achieved despite the targets being predicted not being directly intended for correlation with the mordred descriptors.
Of note is that even though fastprop is approaching the leading performance on this benchmark, other performance metrics cast doubt on the model performance. The weighted mean absolute percentage error (wMAPE) on a per-task basis is shown in Table LABEL:qm8_results_table.
Metric | wMAPE |
E1-CC2 | 3.6% |
E2-CC2 | 3.3% |
f1-CC2 | 83.0% |
f2-CC2 | 86.7% |
E1-PBE0 | 3.6% |
E2-PBE0 | 3.3% |
f1-PBE0 | 80.6% |
F2-PBE0 | 86.4% |
E1-CAM | 3.4% |
E2-CAM | 3.0% |
f1-CAM | 77.3% |
f2-CAM | 81.2% |
Average | 43.0% |
At each level of theory (CC2, PBE0, and CAM) fastprop is reaching the limit of chemical accuracy on excitation energies (E1 and E2) but is significantly less accurate on oscillator strengths (f1 and f2). This can at least partially be attributed to dataset itself. Manual analysis reveals that nearly 90% of the molecules in the dataset fall within only 10% of the total range of f1 values, which is highly imbalanced. Additionally that 90% of molecules actual f1 values are all near-zero or zero, which are intentionally less represented in the wMAPE metric. Future literature studies should take this observation into account and perhaps move away from this splitting approach toward one which accounts for this imbalance.
ESOL
First described in 2004 [32] and has since become a critically important benchmark for QSPR/molecular property prediction studies. The dataset includes molecular structure for approximately 1.1k simple organic molecules are their corresponding experimentally measured free energy of solvation. This property is a classic target of QSPR studies and is especially well suited for fastprop.
The CMPNN [8] model, a derivative of Chemprop, is used for comparison. The performance values are not those from the original paper but instead the corrected results shown on the CMPNN GitHub page [33]. These numbers are the average and standard deviation across 5 repetitions, each including 5-fold cross validation (60/20/20 random split), for a total of 25 models. fastprop performance is reported using the same split sizes across 8 repetitions without cross validation; increasing the number of repetitions further did not meaningfully change model performance.
fastprop achieves an RMSE of 0.643 0.048 trailing the CMPNN at 0.547 0.011 but matching Chemprop at 0.665 0.052. The same pattern from previous benchmarks is repeated - fastprop matches the performance of generic learned representation approaches but is outperformed by bespoke modeling techniques. In this case the CMPNN has been designed to perform better on these benchmark datasets specifically and at an increased cost in both execution time and complexity.
FreeSolv
The Free Energy of Solvation (FreeSolv) database is a more curated alternative to ESOL developed and maintained by Mobley et al. in 2014 [34]. It contains approximately 0.6k molecules and their experimental and calculated hydration free energy. This benchmark is the smallest ‘recognized’ benchmark often reported in the molecular property prediction literature.
Again the CMPNN study is used as a reference and the same procedure described in ESOL is followed. fastprop achieves an RMSE of 1.33 0.21, once again trailing trailing the CMPNN at 0.82 0.15 but matching Chemprop at 1.26 0.11.
@@
Using a complex multi-model ensembling method, the reference study achieved an RMSE of 13.2, an MAE of 8.4, and an MAPE of 2.5%. fastprop matches this performance, achieving 13.3 2.1 RMSE, 9.4 0.8 MAE, and 2.6% 0.1% MAPE. Chemprop, however, struggles to match the accuracy of either method. It manages an RMSE of 21.2 2.2 and an MAE of 13.8 2.1 and does not report MAPE.
@@
Assembled by Das and coauthors [29] from a collection of other smaller datasets, this dataset maps ~0.4k molecular structures to a unified-scale Yield Sooting Index (YSI), a molecular property of interest to the combustion community. The reference study performs leave-one-out cross validation to fit a per-fragment contribution model, effectively a training size of >99%, without a holdout set. Though this is not standard practice and can lead to overly optimistic reported performance, the results will be carried forward regardless. The original study did not report overall performance metrics, so they have been re-calculated for this study using the predictions made by the reference model as provided on GitHub 333Predictions are available at this permalink to the CSV file on GitHub.. For comparison fastprop and Chemprop use a more typical 60/20/20 random split and 8 repetitions. Results are summarized in Table LABEL:ysi_results_table.
@@
fastprop significantly outperforms both other models when considering MAE, especially impressive in the case of the reference model which was trained on far more data. When considering RMSE, which penalizes large errors more than MAE, all models are similarly performant. Finally the WMAPE shows that the reference model makes much smaller errors on the highest YSI molecules compared to fastprop. Taken in combination with the MAE and RMSE values, which are respectively worse and competitive with fastprop, the model is likely highly overfit to the training data due to the cross-validation strategy.
Also notable is the difference in training times. Chemprop takes 7 minutes and 2 seconds while fastprop completes in only 38 seconds, again a factor of ten faster.
HOPV15 Subset
The HOPV15 Subset is a collection of ~0.3k organic photovoltaic compounds curated by Eibeck and coworkers from the larger Harvard Organic Photovoltaic (HOPV15 [35]) based on criteria described in their paper [30]. This dataset is unique in that the target property Power Conversion Efficiency is both experimentally measurable or can be derived from quantum mechanics simulations, but regardless is not a ‘classical’ target of QSPR. After applying a variety of established modeling techniques Eibeck et al. achieved a best-case MAE of 1.32 0.10 averaged across 5 randomly selected 60/20/20 data splits by using a simple molecular fingerprint representation and support vector regression.
In the course of this study it was found that changing the random seed when using only 5 repetitions would lead to dramatically different model performance. Thus, for this benchmark the number of repetitions was set higher than the reference study at 15. fastprop reports an average MAE of 1.55 0.20 and an RMSE of 1.93 0.20, in line with the performance of Chemprop at an MAE of 1.60 0.15 and an RMSE of 1.97 0.16. Execution time is again dramatically different with Chemprop taking 11 minutes and 35 seconds whereas fastprop took only 2 minutes and 3 seconds of which 1 minute and 47 seconds was spent calculating descriptors for the abnormally large molecules in HOPV15.
Fubrain
First described by Esaki and coauthors, the Fraction of Unbound Drug in the Brain (Fubrain) dataset is a collection of about 0.3k small molecule drugs and their corresponding experimental experimentally measured unbound fraction in the brain, a critical metric for drug development [22]. This specific target in combination with this dataset size makes this benchmark highly relevant for typical QSPR studies.
The study that first generated this dataset used mordred descriptors but as is convention they strictly applied linear modeling methods. Using both cross validation and and external test sets, they had an effective training/validation/testing split of 0.64/0.07/0.28 which will be repeated 4 times here for comparison. All told, their model achieved an RMSE of 0.53 averaged across all testing data.
In only 44 seconds, of which 36 are spent calculating descriptors, fastprop far exceeds the reference model with an RMSE of 0.19 0.03 Under the same conditions Chemprop approaches fastprop’s performance with an RMSE of 0.22 0.04 but requires 5 minutes and 11 seconds to do so, in this case a 7 times performance improvement for fastprop over Chemprop.
Delta-Fubrain
Also noteworthy for the Fubrain dataset is that it has been subject to the delta-learning approach to small dataset limitations. DeepDelta [14] performed a 90/0/10 cross-validation study of the Fubrain dataset in which the training and testing molecules were used to generate all possible pairs and then the differences in the property 444Although the original Fubrain study reported untransformed fractions, the DeepDelta authors confirmed via GitHub that DeepDelta was trained on log base-10 transformed fraction values, which is replicated here. were predicted rather than absolute values. They reported an RMSE of 0.830 0.023, whereas a Chemprop model trained to directly predict property values was only able to reach an accuracy of 0.965 0.019 when evaluated on its capacity to predict property differences.
fastprop is able to overcome these limitations. Using the same model from above (re-trained to predict log-transformed values), which has less training data than DeepDelta even before the augmentation, fastprop achieves 0.740 0.087 RMSE after pairing all withheld test molecules. Increasing the amount of training data while retaining some samples for early stopping yields only small improvements, showing that fastprop may be approaching the irreducible error of Fubrain. With an 89/1/10 split the RMSE of fastprop decreases to 0.7118 0.1381, though with significantly increased variance due to small size of the testing data. Regardless, the execution time and scaling issues of DeepDelta and the inaccuracy of Chemprop are effectively circumvented by fastprop.
PAHs
@@
For comparison, fastprop and Chemprop are trained using 8 repetitions of a typical 80/10/10 random split - only 44 molecules in the training data. fastprop matches the performance of the bespoke descriptors with a correlation coefficient of 0.976 0.027. This corresponds to an MAE of 0.160 0.035 and an MAPE of 2.229 0.061%. Chemprop effectively fails on this dataset, achieving a correlation coefficient of only 0.59 0.24, an MAE of 1.04 0.33 (one anti-correlated outlier replicate removed). Despite the large parameter size of the fastprop model relative to the training data, it readily outperforms Chemprop in the small-data limit.

### v2

@@ Abstract
Quantitative Structure-Property/Activity Relationship studies, often referred to interchangeably as QS(P/A)R, seek to establish a mapping between molecular structure and an arbitrary Quantity of Interest (QOI). Since its inception this was done on a QOI-by-QOI basis with new descriptors being devised by researchers to specifically map to their QOI. This continued for years and culminated in packages like DRAGON (later E-dragon), PaDEL-descriptor (and padelpy), Mordred, and many others. The sheer number of different packages resulted in the creation of ‘meta-packages’ which served only to aggregate these other calculators, including tools like molfeat, ChemDes, Parameter Client, and AIMSim.
Generalizable descriptor-based modeling was a natural evolution of these meta-packages’ development, though it historically focused almost exclusively on linear methods. Efforts to incorporate Deep Learning (DL) as a regression technique (Deep-QSPR), which would be capable of capturing the non-linear behavior of arbitrary QOIs, have instead focused on using molecular fingerprints as inputs. The combination of bulk molecular-level descriptors with DL has remained largely unexplored, in significant part due to the orthogonal domain expertise required to combine the two. Generalizable QSPR has turned to learned representations primarily via message passing graph neural networks. This approach has proved remarkably effective but is not without drawbacks. Learning a representation requires large datasets to avoid over-fitting or even learn at all, loses interpretability since an embedding’s meaning can only be induced retroactively, and needs significant execution time given the complexity of the underlying message passing algorithm. This paper introduces fastprop, a software package and general Deep-QSPR framework that combines a cogent set of molecular descriptors with DL to achieve state-of-the-art performance on datasets ranging from tens to tens of thousands of molecules. fastprop is designed with Research Software Engineering best practices and is free and open source, hosted at github.com/jacksonburns/fastprop.
@@ Introduction
While iterations on LRs and novel approaches to low-data regimes have been in development, the classical QSPR community has continued their work. A turning point in this domain was the release of mordred, a fast and well-developed package capable of calculating more than 1600 molecular descriptors. Critically this package was fully open source and written in Python, allowing it to readily interoperate with the world-class Python DL software ecosystem that greatly benefitted the LR community. Now despite previous evidence that molecular descriptors cannot achieve generalizable QSPR in combination with DL, the opposite is shown with fastprop.
@@
At its core the fastprop ‘architecture’ is simply the mordred molecular descriptor calculator 222The original mordred package is no longer maintained. fastprop uses a fork of mordred called mordredcommunity that is maintained by community-contributed patches (see github.com/JacksonBurns/mordred-community). Multiple descriptor calculators from the very thorough review by McGibbon et al. [16] could be used instead, though none are as readily interoperable as mordred. [17] connected to a Feedforward Neural Network (FNN) implemented in PyTorch Lightning [18] (Figure 1) - an existing approach formalized into an easy to use, reliable, and correct implementation. The user simply specifies a set of SMILES [19], a linear textual encoding of molecules, and their corresponding properties. fastprop automatically calculates and caches the corresponding molecular descriptors with mordred, re-scales both the descriptors and the targets appropriately, and then trains an FNN with to predict the indicated target properties. By default this FNN is two hidden layers with 1800 neurons each connected by ReLU activation functions, though the configuration can be readily changed via the command line interface or configuration file. fastprop owes its success to the cogent set of descriptors assembled by the developers of mordred, the ease of training FNNs with modern software like PyTorch Lightning, and the careful application of Research Software Engineering best practices that make it as user friendly as the best-maintained alternatives.
@@
# pah.yml # generic args output_directory: pah random_seed: 55 problem_type: regression # featurization input_file: pah/arockiaraj_pah_data.csv target_columns: log_p smiles_column: smiles descriptor_set: all # preprocessing clamp_input: True # training number_repeats: 4 number_epochs: 100 batch_size: 64 patience: 15 train_size: 0.8 val_size: 0.1 test_size: 0.1 sampler: random
@@
There are a number of established molecular property prediction benchmarks commonly used in LR studies, especially those standardized by MoleculeNet [24]. Principal among them are QM8 [25] and QM9 [26], often regarded as the standard benchmark for property prediction. These are important benchmarks and QM9 is included for completeness, though the enormous size and rich coverage of chemical space (inherent in the design of the combinatorial datasets) means that nearly all model architectures are highly accurate, including fastprop.
Real world datasets, particularly those common in QSPR studies, often number in the hundreds. To demonstrate the applicability of fastprop to these regimes, many smaller datasets are selected including some from the QSPR literature that are not established benchmarks. These studies relied on more complex and slow modeling techniques (ARA) or the design of a bespoke descriptor (PAHs) and have not yet come to rely on learned representations as a go-to tool. In these data-limited regimes where LRs sometimes struggle, fastprop and its intuition-loaded initialization are highly powerful. To emphasize this point further, the benchmarks are presented in order of size, descending.
Two additional benchmarks showing the limitations of fastprop are included after the main group of benchmarks: Fubrain and QuantumScents. The former demonstrates how fastprop can outperform LRs but still trail approaches like delta learning. The later is a negative result showing how fastprop can fail on especially difficult, atypical targets.
@@
The method for splitting data into training, validation, and testing sets varies on a per-study basis and is described in each sub-section. Sampling is performed using the astartes package [27] which implements a variety of sampling algorithms and is highly reproducible. For datasets containing missing target values or invalid SMILES strings, those entries were dropped.
Results for fastprop are reported as the average value of a metric and its standard deviation across a number of repetitions (repeated re-sampling of the dataset). The number of repetitions is chosen to either match referenced literature studies or else increased from two until the performance no longer meaningfully changes. Note that this is not the same as cross-validation. Each section also includes the performance of a zero-layer (i.e. linear regression) network as a baseline to demonstrate the importance of non-linearity in a deep NN.
For performance metrics retrieved from literature it is assumed that the authors optimized their respective models to achieve the best possible results; therefore, fastprop metrics are reported after model optimization using the fastprop train ... --optimize option. When results are generated for this study using Chemprop, the default settings are used except that the number of epochs is increased to allow the model to converge and batch size is increased to match dataset size and speed up training. Chemprop was chosen as a point of comparison throughout this study since it is among the most accessible and well-maintained software packages for molecular machine learning.
When reported, execution time is as given by the unix time command using Chemprop version 1.6.1 on Python 3.8 and includes the complete invocation of Chemprop, i.e. time chemprop_train .... The insignificant time spent manually collating Chemprop results (Chemprop does not natively support repetitions) is excluded. fastprop is run on version 1.0.6 using Python 3.11 and timing values are reported according to its internal time measurement which was verified to be nearly identical to the Unix time command. The coarse comparison of the two packages is intended to emphasize the scaling of LRs and Deep-QSPR and that fastprop is, generally speaking, much faster. All models trained for this study were run on a Dell Precision series laptop with an NVIDIA Quadro RTX 4000 GPU and Intel Xeon E-2286M CPU.
@@
Benchmark Results
See Table LABEL:results_table for a summary of all the results. Subsequent sections explore each in greater detail.
@@
p |
QM9 | ~134 | MAE | 0.0047a | 0.0060 | 0.0081a | ~ |
Pgp | ~1.3 | AUROC | 0.94b | 0.90 | 0.89b | ~ |
ARA | ~0.8 | Accuracy | 91c | 0.88 | 82* | 0.083 |
Flash | ~0.6 | RMSE | 13.2d | 13.0 | 21.2* | 0.021 |
YSI | ~0.4 | MAE | 22.3e | 25.0 | 28.9* | 0.29 |
PAH | ~0.06 | R2 | 0.96f | 0.97 | 0.59* | 0.0012 |
a [9] b [28] c [29] d [30] e [31] f [4] * These reference results were generated for this study.
Statistical comparisons of fastprop to Chemprop (shown in the p column) are performed using the non-parametric Wilcoxon-Mann-Whitney Test as implemented in GNumeric. Values are only shown for results generated in this study which are known to be performed using the same methods. Only the results for Flash and PAH are statistically significant at 95% confidence (p<0.05).
@@
fastprop achieves 0.0060 0.0002 mean absolute error, whereas Chemprop achieves 0.00814 0.00001 and the UniMol framework manages 0.00467 0.00004. This places the fastprop framework ahead of previous learned representation approaches but still trailing UniMol. This is not completely unexpected since UniMol encodes 3D information from the dataset whereas Chemprop and fastprop use only 2D. Future work could evaluate the use of 3D-based descriptors to improve fastprop performance in the same manner that UniMol has with LRs. All methods are better than a purely linear model trained on the molecular descriptors, which manages only a 0.0095 0.0006 MAE.
Pgp
The model in the original study uses a molecular interaction field but has since been surpassed by other models. According to TDC the current leader [28] on this benchmark has achieved an AUROC of 0.938 0.002 333See the TDC Pgp leaderboard.. On the same leaderboard Chemprop [6] achieves 0.886 0.016 with the inclusion of additional molecular features. fastprop yet again approaches the performance of the leading methods and outperforms Chemprop, here with an AUROC of 0.903 0.033 and an accuracy of 83.6 4.6%. Remarkably, the linear model outperforms both Chemprop and fastprop, approaching the performance of the current leader with an AUROC of 0.917 0.016 and an accuracy of 83.8 3.9%.
ARA
Compiled by Schaduangrat et al. in 2023 [29], this dataset maps ~0.8k small molecules to a binary label indicating if the molecule is an Androgen Receptor Antagonist (ARA). The reference study introduced DeepAR, a highly complex modeling approach, which achieved an accuracy of 0.911 and an AUROC of 0.945.
For this study an 80/10/10 random splitting is repeated four times on the dataset since no analogous split to the reference study can be determined. Chemprop takes 16 minutes and 55 seconds to run on this dataset and achieves only 0.824 0.020 accuracy and 0.898 0.022 AUROC. fastprop takes only 1 minute and 54 seconds (1 minute and 39 seconds for descriptor calculation) and is competitive with the reference study in performance, achieving a 88.2 3.7% accuracy and 0.935 0.034 AUROC. The purely model falls far behind these methods with a 71.8 6.6% accuracy and 0.824 0.052 AUROC, as expected.
@@
Using a complex multi-model ensembling method, the reference study achieved an RMSE of 13.2, an MAE of 8.4, and an MAPE of 2.5%. fastprop matches this performance, achieving 13.0 2.0 RMSE, 9.0 0.5 MAE, and 2.7% 0.1% MAPE. Chemprop, however, struggles to match the accuracy of either method. It manages an RMSE of 21.2 2.2 and an MAE of 13.8 2.1 and does not report MAPE. This is worse than the performance of the linear model, with an RMSE of 16.1 4.0, an MAE of 11.3 2.9, and an MAPE of 3.36 0.77%
@@
Assembled by Das and coauthors [31] from a collection of other smaller datasets, this dataset maps ~0.4k molecular structures to a unified-scale Yield Sooting Index (YSI), a molecular property of interest to the combustion community. The reference study performs leave-one-out cross validation to fit a per-fragment contribution model, effectively a training size of >99%, without a holdout set. Though this is not standard practice and can lead to overly optimistic reported performance, the results will be carried forward regardless. The original study did not report overall performance metrics, so they have been re-calculated for this study using the predictions made by the reference model as provided on GitHub 444Predictions are available at this permalink to the CSV file on GitHub.. For comparison fastprop and Chemprop use a more typical 60/20/20 random split and 8 repetitions. Results are summarized in Table LABEL:ysi_results_table.
@@
Linear | 82 39 | 180 120 | 0.47 0.23 |
fastprop again outperforms Chemprop, in this case approaching the overly-optimistic performance of the reference model. Taking into account that reference model has been trained on a significantly larger amount of data, this performance is admirable. Also notable is the difference in training times. Chemprop takes 7 minutes and 2 seconds while fastprop completes in only 42 seconds, again a factor of ten faster. The linear model fails entirely, performing dramatically worse than all other models.
PAH
@@
For comparison, fastprop and Chemprop are trained using 8 repetitions of a typical 80/10/10 random split - only 44 molecules in the training data. fastprop matches the performance of the bespoke descriptors with a correlation coefficient of 0.972 0.025. This corresponds to an MAE of 0.19 0.10 and an MAPE of 2.5 1.5%. Chemprop effectively fails on this dataset, achieving a correlation coefficient of only 0.59 0.24, an MAE of 1.04 0.33 (one anti-correlated outlier replicate removed). This is worse even than the purely linear model, which manages a correlation coefficient of 0.78 0.22, an MAE of 0.59 0.22, and an RMSE of 0.75 0.32. Despite the large parameter size of the fastprop model relative to the training data, it readily outperforms Chemprop in the small-data limit.